The nucleus of 8Be, which consists of 4 protons and 4 neutrons, is very unstable and spontaneously breaks into two alpha particles (helium nuclei, each consisting of 2 protons and 2 neutrons).

(a) What is the force between the two alpha particles when they are 3.82 10−15 m apart?
N

(b) What is the initial magnitude of the acceleration of the alpha particles due to this force? Note that the mass of an alpha particle is 4.0026 u.
m/s2
Answer #1

Solution)

Part a) Force F= kq^2/r^2

F=(9*10^9)*(2*1.6*10^-19)^2/(3.82*10^-15)^2

F= 63.15 N

=====

Part b) acceleration, a= F/m =63.15/(4.0026*1.66*10^-27)

=9.5*10^27 m/s^2 (Ans)

======

Good luck!:)